The number midway between -6 and 2 is ________?
I'm not confident that I know how to solve this the right way, please help if you can.
Thank you!

Answer by RAY100(1637)

Midpoint is (X1+X2)/2
.
in this case {-6+2}/2 =-4/2=-2
.
check,, distance from -6 to -2 is 4,,,,,and distance from +2 to-2 is also 4
.
.
.
Lets use a simple example ,,,,4,5,6,,,obviously 5 is midpoint,
.
to use math,,we could do,,,6-4 =2,,,and 2/2=1,,and then 6-1=5,,or 4+1=5
.
but it is easier to do (4+6) = 10,,,,10/2=5
.